# Memory-Enhanced Neural Solvers for Efficient Adaptation in Combinatorial Optimization

### v1

@@ 4 Experiments
We evaluate our method across widely recognized CO problems Travelling Salesman (TSP) and Capacitated Vehicle Routing (CVRP). These problems serve as standard benchmarks for evaluating RL-based CO methods (Deudon et al., 2018; Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023; Hottung et al., 2022; Chalumeau et al., 2023b). In Section 4.1, we validate our method by comparing it to leading approach for single-policy adaptation using online collected data, EAS (Hottung et al., 2022). We present their comparative performance in line with established standard benchmarking from literature, and elucidate the mechanisms employed to adapt the action distribution of the base policy during inference. We then demonstrate how MEMENTO can be combined with SOTA neural solver COMPASS with no additional retraining in Section 4.2. Finally, we scale those methods to larger instances in Section 4.3, outperforming heatmap-based methods.
@@
We provide access to the code [^1] Code, checkpoints, and evaluation sets available at https://github.com/instadeepai/memento utilized for training our method and executing all baseline models. Additionally, we release our checkpoints for all problem types and scales, accompanied by the necessary datasets to replicate our findings. We implement our method and experiments in JAX (Bradbury et al., 2018), in a codebase that is compatible with previous JAX implementation released in the community. The two problems are also JAX implementations from Jumanji (Bonnet et al., 2023), enabling us to fully leverage hardware accelerators such as TPUs. Our implementations are optimized for TPU v3-8, aligning with the hardware utilized in our experiments.
Figure 2: Performance improvement of MEMENTO and EAS over a base policy. Results are % relative to a greedy rollout of the given base policy (POMO). The benchmark comprises $8$ datasets with $4$ distinct instance sizes on both TSP (left) and CVRP (right). MEMENTO is consistently outperforming policy-gradient based method EAS, with significant improvement to the base policy.
@@ Results
The average performance of each method across all problem settings are presented in Fig. 2. The observations we draw are three-fold. First, MEMENTO outperforms its base model (POMO) on the entire benchmark by a significant margin: showing that its adaptive search is superior to stochastic sampling. Second, MEMENTO outperforms EAS on all $8$ tasks (spanning both in- and out-of-distribution) for both environments, highlighting the efficacy of meta-learned policy updates when compared to vanilla policy gradients. Finally, we note that this improvement is significant; for example, on TSP100, MEMENTO is doing $60$% better than sampling, while EAS only $6$%.
To better understand how MEMENTO uses its memory to adapt the action distribution of its base policy, we analyse the logit update that an action gets with respect to its associated data, akin to similar analysis conducted in meta-learning (Lu et al., 2022; Jackson et al., 2024), and compare it to the REINFORCE policy gradient. On Fig. 3, we plot the heatmap of the logit update with respect to the log probability (logp) and return associated with an action.
We observe that the main rules learned by MEMENTO are similar to REINFORCE: an action with low logp and high return gets a positive update while an action with high logp and low return gets discouraged. It is nonetheless interesting to see discrepancies. In particular, MEMENTO’s rules are not symmetric with respect to $x=0$, it only encourages action that are above the mean return, but gives higher amplitude to high-return actions than it discourages low-return ones, pushing for performance while preserving exploration. Note that we also expect some kind of dissymetry from EAS since it combines REINFORCE with a term that increases the likelihood to generate the best solution ever found. Finally, the heatmap shows that the update amplitude provided by MEMENTO is significantly bigger than REINFORCE, explaining the higher data-efficiency observed.
4.2 Zero-shot combination of MEMENTO with new solvers
@@
We load COMPASS checkpoint from Chalumeau et al. (2023b), and we reuse the checkpoint of MEMENTO that was trained using POMO and combine it with COMPASS. For the first half the budget, we do not use MEMENTO, to let COMPASS find the appropriate pre-trained policy. Once the search starts narrowing down, we activate MEMENTO to adapt the pre-trained policy, with no need to turn off COMPASS’ search. Fig. 4 shows the typical trends observed at inference. We see the typical search reported in Chalumeau et al. (2023b). At $50$% of budget consumption, MEMENTO is activated and one can observe a clear and significant drop in averaged tour length of the latest sampled solutions. The standard deviation also illustrates how the search is suddenly narrowed down.
@@ 4.3 Scaling RL construction methods to larger instances
Scaling neural solvers to larger instances is a crucial challenge to be tackled in order to bridge the gap to industrial applications. Construction-based solvers are promising approaches, requiring limited expert knowledge whilst providing strong performance. But are still hardly ever trained on larger scales with RL. As a result, their performance reported in the literature is usually quite unfair, being evaluated in large scales instances (>$500$) although having been trained on $100$ nodes. In this section, we explain how we have trained construction solvers POMO and COMPASS on instances of size $500$ with RL (for both TSP and CVRP, checkpoints to be open-sourced); and then use them as new base models to validate properties of MEMENTO at that scale.
@@
Training POMO on instances of size $500$ involves three main steps. First, inspired by curriculum-based methods, we start from a checkpoint pre-trained on TSP100. Second, we use Efficient Attention (Rabe and Staats, 2022) to reduce the memory cost of multi-head attention, enabling to rollout problems in parallel despite the $O(n^{2})$ memory requirement. Third, we use gradient accumulation to keep good estimates despite the constrained smaller instance batch size. These three combined tricks enable to train POMO till convergence within $4$ days, and consequently COMPASS and MEMENTO. We also build the zero-shot combined MEMENTO (COMPASS) checkpoint. We release all checkpoints (TSP & CVRP) publicly to ease future research at this scale.
Figure 5: Performance improvement of MEMENTO over EAS on instances of size $500$ on TSP (left) and CVRP (right) for increasing number of sequential attempts and sizes of attempt batch. Green areas correspond to setting for which MEMENTO’s online adaptation is more efficient than EAS. MEMENTO always outperforms EAS on TSP, and for most settings on CVRP. The difference is particularly significant for low budgets, showing the data-efficiency and robustness of MEMENTO.
@@
On those larger instances, the cost of constructing a solution is significantly increased (for the architecture considered: squared in computation and memory cost, and linear in number of sequential operations). It is also harder to get numerous parallel shots, which most methods depend on. It hence becomes crucial to develop data-efficient methods which stay robust to low data regimes. To assess MEMENTO’s efficiency, we train it on instances of size $500$ (for both TSP and CVRP, checkpoints released) and compare its adaptation to EAS on a set of unseen instances, for a range of budget expressed in number of sequential attempts and size of batched solutions per attempts. Figure 5 reports the percentage of improvement (in absolute cost) brought by MEMENTO over EAS. From this figure, we can observe three main tendencies. First, the data-efficiency of MEMENTO is illustrated by its consistent superiority on low budget (lower left of plots), reaching significant improvement compared to EAS. Then, we can see that, on CVRP, the gap increases with the sequential budget for low parallelism. It demonstrates that the meta-learned rule of MEMENTO stays robust, whereas the gradient estimates of EAS get deteriorated and cannot bring further improvement despite additional sequential budget. To finish with, we can see that MEMENTO fully dominates the heatmap on TSP, even for higher budget. Nevertheless, when budget becomes bigger, EAS is able to outperform MEMENTO, under the condition that batches are large enough.
Table 2: Results of MEMENTO against the baseline algorithms on instances of size $500$ of TSP (a) and CVRP (b) for two budget regimes (low and high). We observe that POMO-based methods are outperforming concurrent methods. MEMENTO achieves single-agent SOTA (resp. overall SOTA) on $3$ out of $4$ settings when combined with POMO (resp. COMPASS).
@@ Results
In Table 2, we report performance of our checkpoints, using datasets introduced in Fu et al. (2021) and commonly used in the literature. For each environment, we report the results on two different setting: (i) Low Budget, where the methods are given 25 0002500025\,00025 000 attempts, (ii) High Budget, where 100 000100000100\,000100 000 attempts are available. We also report results from concurrent RL methods (Qiu et al., 2022; Dernedde et al., 2024; Sun and Yang, 2023; Ye et al., 2023), without method-agnostic local search; and industrial solvers LKH (Helsgaun, 2017) and Concorde (Applegate et al., 2006). Interestingly, stochastically sampling solutions with POMO for less than $10$ minutes already provides competitive results, ranking second among the five leading neural solvers. Adding EAS on top of POMO enables to compete with leading method MOCO on TSP500.
It is the first time that POMO is trained on those instance sizes, making it the first RL-trained neural solver that can be used for large instances both on TSP and CVRP, since most previous large scale RL methods are graph-specific, and hence cannot be applied on CVRP. On this benchmark, MEMENTO achieved state-of-the-art on $3$ of the $4$ tasks when combined with COMPASS, showing significant improvement compared to previous state-of-the-art MOCO. It also proves efficiency on those large scales by being the best single agent method on all TSP instances, and on the low budget CVRP task.
@@ 5 Conclusion
We present MEMENTO, a method to improve adaptation of neural CO solvers to unseen instances. MEMENTO enables to condition a policy directly on data collected online during search, facilitating rapid adaptation and efficient use of the search budget. In practice, it proves to outperform stochastic sampling and policy-gradient fine-tuning of the policy, and shows it can be train with one solver, and combined with other existing solvers with no-retraining needed.
We demonstrate the efficacy of MEMENTO by achieving state-of-the-art single-policy adaptation on the standard benchmark on TSP and CVRP, both in and out-of-distribution. Moreover, we find that MEMENTO realises interpretable update rules to the underlying policy; trading off exploration and exploitation over the search budget and rediscovering REINFORCE-style updates. We further demonstrate the flexibility of this approach by achieving zero-shot combination of MEMENTO and COMPASS, achieving state-of-the-art on a benchmark of $12$ tasks. The simplicity and efficacy of this method makes it amenable to a very large set of problems, laying another stone towards the application of neural solvers to real-world problems.
Limitations and future work. MEMENTO incurs additional compute processing and memory-usage when compared to the memory-less base policies which it augments. Practically, we find that in all considered settings, the performance gain outweighs these overheads, however practitioners should ensure that this remains the case in other settings. The adaptation enabled by MEMENTO is determined by the data collected by the base policy, therefore this could limit the efficiency when evaluation instances are very far from the training distribution, or otherwise when the base policy has harmfully poor initial performance. Finally, we believe that the capacity of MEMENTO could be improved if trained with broader distribution, in particular using variable instance sizes.
Acknowledgements
Research supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
@@ A.1 Standard benchmark
We evaluate our method, MEMENTO, against leading RL methods and industrial solvers. For RL specific methods, we provide results for POMO with greedy action selection and stochastic sampling, and EAS; an active search RL method built on POMO, that fine-tunes a policy on each problem instance. We also report population-based method COMPASS and the zero-shot combination of COMPASS with MEMENTO.
@@
We compare to the heuristic solver LKH3 [Helsgaun, 2017]; the current leading industrial solver of both TSP and CVRP, as well as an exact solver Concorde [Applegate et al., 2006] which is a TSP-specific industrial solver.
@@
We use datasets of 10,000 instances with 100 cities/customer nodes drawn from the training distribution, and three generalization datasets of 1,000 instances of sizes 125, 150, and 200, all from benchmark sets frequently used in the literature [Kool et al., 2019, Kwon et al., 2020, Hottung et al., 2022, Grinsztajn et al., 2023, Chalumeau et al., 2023b]. Tables  3(a) and  3(b) display results for TSP and CVRP on the standard benchmark, respectively. The columns provide the absolute tour length, the optimality gap, and the total inference time for each method.
@@
We can see that in the single-agent setting, MEMENTO leads the entire benchmark, and in the population-based setting, MEMENTO (COMPASS) is leading the whole benchmark. MEMENTO is able to give significant improvement to COMPASS on CVRP, pushing significantly the state-of-the-art on this benchmark. On TSP, the improvement is not significant enough to be visible on the rounded results. To finish with, we can observe that the time cost associated with MEMENTO is reasonable and is worth the performance improvement (except for MEMENTO (COMPASS)).
Table 3: Results of MEMENTO against the baseline algorithms for (a) TSP, (b) CVRP. The methods are evaluated on instances from training distribution as well as on larger instance sizes to test generalization.
@@
Concorde 7.765 $0.000\%$ 82M 8.583 $0.000\%$ 12M 9.346 $0.000\%$ 17M 10.687 $0.000\%$ 31M
LKH3 7.765 $0.000\%$ 8H 8.583 $0.000\%$ 73M 9.346 $0.000\%$ 99M 10.687 $0.000\%$ 3H
POMO (greedy) 7.796 0.404% 16S 8.635 0.607% 20S 9.440 1.001% 29S 10.933 2.300% 45S

### v2

@@ 4 Experiments
We evaluate our method across widely recognized CO problems Travelling Salesman (TSP) and Capacitated Vehicle Routing (CVRP). These problems serve as standard benchmarks for evaluating RL-based CO methods (Deudon et al., 2018; Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023; Hottung et al., 2022; Choo et al., 2022; Chalumeau et al., 2023b). In Section 4.1, we validate our method by comparing it to leading approach for single-policy adaptation using online collected data, EAS (Hottung et al., 2022). We present their comparative performance in line with established standard benchmarking from literature, and elucidate the mechanisms employed to adapt the action distribution of the base policy during inference. We then demonstrate how MEMENTO can be combined with SOTA neural solver COMPASS with no additional retraining in Section 4.2. Finally, we scale those methods to larger instances in Section 4.3, outperforming heatmap-based methods.
(a) TSP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Obj. Gap Obj. Gap Obj. Gap
LKH3 7.765 $0.000\%$ 8.583 $0.000\%$ 9.346 $0.000\%$ 10.687 $0.000\%$
POMO (greedy) 7.796 0.404% 8.635 0.607% 9.440 1.001% 10.933 2.300%
POMO (sampling) 7.779 0.185% 8.609 0.299% 9.401 0.585% 10.956 2.513%
SGBS* 7.769* $0.058\%$ 9.367* $0.220\%$ 10.753* 0.619%
EAS-Emb 7.778 0.161% 8.604 0.238% 9.380 0.363% 10.759 0.672%
MEMENTO 7.768 0.045% 8.592 0.109% 9.365 0.202% 10.758 0.664%
(b) CVRP
Training distr. Generalization
$n=100$ $n=125$ $n=150$ $n=200$
Method Obj. Gap Obj. Gap Obj. Gap Obj. Gap
LKH3 15.65 $0.000\%$ 17.50 $0.000\%$ 19.22 $0.000\%$ 22.00 $0.000\%$
POMO (greedy) 15.874 1.430% 17.818 1.818% 19.750 2.757% 23.318 5.992%
POMO (sampling) 15.713 0.399% 17.612 0.642% 19.488 1.393% 23.378 6.264%
SGBS* 15.659* $0.08\%$ 19.426* $1.08\%$ 22.567* $2.59\%$
EAS-Emb 15.663 0.081% 17.536 0.146% 19.321 0.528% 22.541 2.460%
MEMENTO 15.657 0.066% 17.521 0.095% 19.317 0.478% 22.492 2.205%
Figure 3: Results of MEMENTO against baselines on a standard benchmark comprising $8$ datasets with $4$ distinct instance sizes on TSP (a) and CVRP (b). Methods are evaluated on instances from training distribution ($100$) as well as on larger instance sizes to test generalization. The right panel aggregates those results, reported in % improvement relative to a greedy rollout of the given base policy (POMO). MEMENTO is consistently outperforming policy-gradient based method EAS and tree-search based method SGBS, with significant improvement over the base policy. *SGBS results are reported from Choo et al. (2022), inducing biases in the comparison, discussed in Section A.1.
@@
We provide access to the code [^1] Code, checkpoints, and evaluation sets at https://github.com/instadeepai/memento utilized for training our method and executing all baseline models. Additionally, we release our checkpoints for all problem types and scales, accompanied by the necessary datasets to replicate our findings. We implement our method and experiments in JAX (Bradbury et al., 2018), in a codebase that is compatible with previous JAX implementation released in the community. The two problems are also JAX implementations from Jumanji (Bonnet et al., 2023), enabling us to fully leverage hardware accelerators such as TPUs. Our implementations are optimized for TPU v3-8, aligning with the hardware utilized in our experiments.
@@ Results
The average performance of each method across all problem settings are presented in Fig. 3. The observations we draw are three-fold. First, MEMENTO outperforms its base model (POMO) on the entire benchmark by a significant margin: showing that its adaptive search is superior to stochastic sampling. Second, MEMENTO outperforms EAS on all $8$ tasks (spanning both in- and out-of-distribution) for both environments, highlighting the efficacy of learned policy updates when compared to vanilla policy gradients. Finally, we note that this improvement is significant; for example, on TSP100, MEMENTO is doing $60$% better than sampling, while EAS only $6$%.
Analysing the update rule
To better understand how MEMENTO uses its memory to adapt the action distribution of its base policy, we analyse the logit update that an action gets with respect to its associated data, akin to similar analysis conducted in meta-learning (Lu et al., 2022; Jackson et al., 2024), and compare it to the REINFORCE policy gradient. On Fig. 4, we plot the heatmap of the logit update with respect to the log probability (logp) and return associated with an action. We observe that the main rules learned by MEMENTO are similar to REINFORCE: an action with low logp and high return gets a positive update while an action with high logp and low return gets discouraged. It is nonetheless interesting to see discrepancies. In particular, MEMENTO’s rules are not symmetric with respect to $x=0$, it only encourages action that are above the mean return, but gives higher amplitude to high-return actions than it discourages low-return ones, pushing for performance while preserving exploration. A similar dissymetry would be expected with EAS since it combines REINFORCE with a term that increases the likelihood to generate the best solution found.
Additionally, MEMENTO uses more inputs than REINFORCE: current budget available, trajectory logp, age of the data, and previous MEMENTO logit are also used to derive the new action logit. Appendix G provides an ablation study of those additional inputs, validating their importance.
4.2 Zero-shot combination of MEMENTO with unseen solvers
@@
We combine our checkpoint of MEMENTO with a released model of COMPASS (Chalumeau et al., 2023b). For the first half of the budget, we do not use MEMENTO, to let COMPASS find the appropriate pre-trained policy. Once the search starts narrowing down, we activate MEMENTO to adapt the pre-trained policy, with no need to turn off COMPASS’ search. Fig. 5 shows the typical trends observed at inference. We see the typical search reported in Chalumeau et al. (2023b). At $50$% of budget consumption, MEMENTO is activated and one can observe a clear and significant drop in averaged tour length of the latest sampled solutions. The standard deviation also illustrates how the search is suddenly narrowed down.
@@ 4.3 Scaling RL construction methods to larger instances
Scaling neural solvers to larger instances is a crucial challenge to be tackled in order to bridge the gap to industrial applications. Auto-regressive construction-based solvers are promising approaches, requiring limited expert knowledge whilst providing strong performance. But are still hardly ever trained on larger scales with RL. As a result, their performance reported in the literature is usually quite unfair, being evaluated in large scales instances (>$500$) although having been trained on $100$ nodes. In this section, we explain how we have trained construction solvers POMO and COMPASS on instances of size $500$ with RL (for both TSP and CVRP, checkpoints to be open-sourced); and then use them as new base models to train and validate properties of MEMENTO at that scale.
Figure 6: Performance improvement of MEMENTO over EAS on instances of size $500$ on TSP (left) and CVRP (right) for increasing number of sequential attempts and sizes of attempt batch. Green areas correspond to settings for which MEMENTO’s online adaptation is more efficient than EAS. MEMENTO always outperforms EAS on TSP, and for most settings on CVRP. The difference is particularly significant for low budgets, showing the data-efficiency and robustness of MEMENTO.
@@
Training POMO on instances of size $500$ involves three main steps. First, inspired by curriculum-based methods, we start from a checkpoint pre-trained on TSP100. Second, we use Efficient Attention (Rabe & Staats, 2022) to reduce the memory cost of multi-head attention, enabling to rollout problems in parallel despite the $O(n^{2})$ memory requirement. Third, we use gradient accumulation to keep good estimates despite the constrained smaller instance batch size. These three combined tricks enable to train POMO till convergence within $4$ days, and consequently to train COMPASS and MEMENTO. We also build the zero-shot combined MEMENTO (COMPASS) checkpoint. We release all checkpoints (TSP & CVRP) publicly to ease future research at this scale.
Inference-time constraints
On those larger instances, the cost of constructing a solution is significantly increased (for POMO’s architecture: squared in computation and memory, and linear in sequential operations). It is also harder to get numerous parallel shots, which most methods depend on. It hence becomes crucial to develop data-efficient methods which stay robust to low budget regimes; and important to know which method to use for each budget constraint.
@@
We compare MEMENTO and EAS on a set of 128 unseen instances (of size $500$), for a range of budget expressed in number of sequential attempts and number of parallel attempts attempts. Figure 6 reports the percentage of improvement (in absolute cost) brought by MEMENTO over EAS. From this figure, we can observe three main tendencies. First, the data-efficiency of MEMENTO is illustrated by its consistent superiority on low budget (lower left of plots), reaching significant improvement compared to EAS. Then, we can see that, on CVRP, the gap increases with the sequential budget for low parallelism. It demonstrates that MEMENTO’s update stays robust, whereas the gradient estimates of EAS get deteriorated and cannot bring further improvement despite additional sequential budget. To finish with, we can see that MEMENTO fully dominates the heatmap on TSP, even for higher budget. Nevertheless, when budget for CVRP becomes bigger, EAS is able to outperform MEMENTO, given that batches are large enough ($>80$).
@@
Figure 7: Tables on the left report results of MEMENTO against the baseline algorithms on instances of size $500$ of TSP (a) and CVRP (b) for two budget regimes (low and high). We observe that POMO-based methods are outperforming concurrent methods. MEMENTO achieves single-agent SOTA (resp. overall SOTA) on $3$ out of $4$ settings when combined with POMO (resp. COMPASS). The plots on the right show the time complexity of MEMENTO and EAS for increasing values of decoder size, and instance size. It demonstrates that MEMENTO scales better in time than EAS.
@@ Results
In Fig. 7, we report performance of our checkpoints, using datasets introduced in Fu et al. (2021) and commonly used in the literature. For each environment, we report the results on two different setting: (i) Low Budget, where the methods are given 25 0002500025\,00025 000 attempts, (ii) High Budget, where 100 000100000100\,000100 000 attempts are available. We also report results from concurrent RL methods (Qiu et al., 2022; Dernedde et al., 2024; Sun & Yang, 2023; Ye et al., 2023), without method-agnostic local search; and industrial solvers LKH3 (Helsgaun, 2017) and Concorde (Applegate et al., 2006).
First, we observe that stochastically sampling solutions with POMO for less than $10$ minutes already provides competitive results, ranking second among the five leading neural solvers. Adding EAS on top of POMO enables to compete with leading method MOCO on TSP500. This shows that auto-regressive RL constructive methods are already competitive at this scale. It is the first time that POMO is trained on those instance sizes, making it the first RL-trained neural solver that can be used for large instances both on TSP and CVRP, since most previous large scale RL methods are graph-specific, and hence cannot be applied on CVRP.
Furthermore, those results confirm that MEMENTO’s adaptation performance scales to instance size, and that it can still be zero-shot combined with COMPASS. On this benchmark, MEMENTO achieved state-of-the-art on $3$ of the $4$ tasks when zero-shot combined with COMPASS, showing significant improvement compared to previous state-of-the-art MOCO: the gap to optimality is reduced by more than $10\%$. It also proves efficiency on those large scales by being the best single agent method on all TSP instances, and on the low budget CVRP task.
Time complexity
Time performance at scale is key for the wide adoption of neural solvers since the number of sequential batches of attempts that can be achieved within a time budget will mostly depend on it (plus on the accessible hardware). Adaptation methods must be able to tackle large instances in reasonable time, and they must also scale well with the size of the base solver used, since Luo et al. (2023) demonstrated the importance of using large and multi-layered decoders for performance in neural CO (although they could not yet train them with RL).
Hence, we evaluate MEMENTO and EAS on a set of increasing instance sizes and decoder sizes and report their time complexity on the right part of Fig. 7. These plot show that, despite being slower for small settings, MEMENTO’s adaptation mechanism scales better than EAS. For an instance size of 1000, MEMENTO becomes $20\%$ faster to use the budget. And even for small instances (size $100$), using $10$ layers in the decoder’s architecture (1M parameters) makes EAS $40\%$ slower than MEMENTO. These scaling laws illustrate another benefit of using an approach that learns the update rather than relying on back-propagation to adapt neural solvers at inference time.
@@ 5 Conclusion
We present MEMENTO, a method to improve adaptation of neural CO solvers to unseen instances. MEMENTO enables to condition a policy directly on data collected online during search, facilitating rapid adaptation and efficient use of the search budget. In practice, it proves to outperform stochastic sampling, tree search and policy-gradient fine-tuning of the policy, and shows it can combined with another solver than the one used during training, with no further re-training needed. Additionally, MEMENTO respects several key properties, like robustness to low-budget regimes, and favorable time performance scalability, both in instance size and base solver size.
We demonstrate the efficacy of MEMENTO by achieving state-of-the-art single-policy adaptation on a standard benchmark on TSP and CVRP, both in and out-of-distribution. Moreover, we find that MEMENTO realises interpretable update rules to the underlying policy; trading off exploration and exploitation over the search budget and outperforming REINFORCE-style updates. We further demonstrate the flexibility of this approach by achieving zero-shot combination of MEMENTO and COMPASS, achieving overall state-of-the-art on a benchmark of $12$ tasks. The simplicity and efficacy of this method makes it amenable to a very large set of problems, laying another stone towards the application of neural solvers to real-world problems.
Limitations and future work. MEMENTO incurs additional compute and memory-usage compared to the memory-less base policies which it augments. Practically, we find that the performance gain significantly outweighs these overheads. A potential mitigation could be to derive the mathematical update learned by MEMENTO to avoid relying on the MLP computation. The adaptation enabled by MEMENTO is determined by the data collected by the base policy, therefore this could limit the efficiency when evaluation instances are very far from the training distribution, or when the base policy has harmfully poor initial performance. Finally, we believe that the capacity of MEMENTO could be improved by training on broader distributions, in particular using variable instance sizes.
@@ A.1 Standard benchmark
We evaluate our method, MEMENTO, against leading RL methods and industrial solvers. For RL specific methods, we provide results for POMO with greedy action selection and stochastic sampling, and EAS; an active search RL method built on top of POMO, that fine-tunes a policy on each problem instance. We also report population-based method COMPASS and the zero-shot combination of COMPASS with MEMENTO.
@@
We compare to the heuristic solver LKH3 (Helsgaun, 2017); the current leading industrial solver of both TSP and CVRP, as well as an exact solver Concorde (Applegate et al., 2006) which is a TSP-specific industrial solver, and the CVRP-specific solver (Vidal et al., 2012).
@@
We use datasets of 10,000 instances with 100 cities/customer nodes drawn from the training distribution, and three generalization datasets of 1,000 instances of sizes 125, 150, and 200, all from benchmark sets frequently used in the literature (Kool et al., 2019; Kwon et al., 2020; Hottung et al., 2022; Grinsztajn et al., 2023; Chalumeau et al., 2023b). Tables  2(a) and  2(b) display results for TSP and CVRP on the standard benchmark, respectively. The columns provide the absolute tour length, the optimality gap, and the total inference time that each method takes to solve one instance within the attempts budget.
Times reported
When possible, we decide to report the time to solve one instance rather the entire dataset following four main observations: (i) First, the literature use datasets of varying sizes, e.g. 10k for TSP100, 1k for TSP200, 128 for TSP500, hence time reported can be confusing, and do not enable to clearly see how methods’ solving time scale with instance size. (ii) Second, the default real-world application consist in solving one instance at a time, or solving several instances at the same time but on separated hardware. (iii) Additionally, measuring on the entire dataset mixes several aspects, i.e. the instance solving time is mixed with the batch scalability of the method. We do think that this property is very interesting to know, but should be considered on the side, rather than mixed with the instance solving time. (iv) Moreover, this property will express differently depending on the hardware available. For instance, on a small hardware POMO sampling with $n$ attempts will be $n$ times slower than POMO greedy; but given a large enough GPU, POMO sampling can be parallelised $n$ times and hence take exactly the same amount of time as POMO greedy.
We were able to do so on the standard benchmark since we have implementations of most methods we were comparing in a single framework. Since we are reporting several external methods in Fig. 7(a), we could only report time taken for the full dataset in that case.
Additional comments about the results
@@
On Table 2, we can see that in the single-agent setting, MEMENTO leads the entire benchmark, and in the population-based setting, MEMENTO (COMPASS) is leading the whole benchmark. MEMENTO is able to give significant improvement to COMPASS on CVRP, pushing significantly the state-of-the-art on this benchmark. On TSP, the improvement is not significant enough to be visible on the rounded results. To finish with, we can observe that the time cost associated with MEMENTO is reasonable and is worth the performance improvement (except maybe for TSP results of MEMENTO (COMPASS)).
Notes concerning SGBS
All neural methods reported in Table 2 are our own runs with standardised checkpoints and rollout strategies, except for the results of SGBS, which were taken from Choo et al. (2022). This introduces three biases: (i) the base POMO checkpoint used by SGBS is not exactly the same as our re-trained POMO checkpoint (ii) SGBS uses the domain-specific augmentation trick that we do not use (iii) SGBS pre-selects starting points in CVRP, which we do not do.
Overall, the results reported in SGBS show that SGBS alone is always outperformed by EAS; hence, it should be expected that in all the settings where we outperform EAS, we would significantly outperform SGBS if both methods were evaluated exactly in the same way. We hence expect the gap between MEMENTO and SGBS to be larger than the one reported here. Additionally, SGBS has yet never been validated on larger instances. Nevertheless, we think that SGBS is a very efficient method from the NCO toolbox and appreciate that MEMENTO and SGBS are orthogonal, and could be combined for further improvement. We leave this for future work.
Table 2: Results of MEMENTO against the baseline algorithms for (a) TSP, (b) CVRP. The methods are evaluated on instances from training distribution ($n=100$) as well as on larger instance sizes to test generalization. We use the same dataset as the rest of the literature, those contain 10 0001000010\,00010 000 instances for $n=100$ and $1000$ instances for $n=125,150,200$. Gaps are computed relative to LKH3. We report time needed to solve one instance. *SGBS results are reported from Choo et al. (2022), and do not use the same POMO checkpoint as other reported results. Additionally, they rely on problem-specific tricks that were not used by other methods. Details in Section A.1.
@@
Concorde 7.765 $0.000\%$ 0.49S 8.583 $0.000\%$ 0.72S 9.346 $0.000\%$ 1S 10.687 $0.000\%$ 1.86S
LKH3 7.765 $0.000\%$ 2.9S 8.583 $0.000\%$ 4.4S 9.346 $0.000\%$ 6S 10.687 $0.000\%$ 11S
POMO (greedy) 7.796 0.404% 0.16S 8.635 0.607% 0.2S 9.440 1.001% 0.29S 10.933 2.300% 0.45S